Solve the equation.
4(2z - 3) = 7(z + 2)

Answer by tutorcecilia(2152)

8z - 12 = 7z + 14 [use the distributive property to solve]
8z-7z-12=7z-7z+14
z-12=14
z-12+12=14+12
z=26
.
check by plugging in (z=26) back into the original equation:
4(2z - 3) = 7(z + 2)
4(2(26) - 3) = 7((26) + 2)
4(49) = 7(28)
196=196